# LOTFormer: Doubly-Stochastic Linear Attention via Low-Rank Optimal Transport

### v1

@@ Abstract
[…]
Keywords Linear Attention, Doubly-Stochastic Attention, Transformers, Optimal Transport
@@ 1 Introduction
On the other hand, prior work has shown that the row-stochastic nature of attention matrices often leads to attention concentrating disproportionately on a few tokens, which can hinder effective information flow (Sander et al., 2022; Shahbazi et al., 2025). In the case of vision Transformers, this phenomenon—referred to as token overfocusing—has been observed to degrade performance, and promoting broader token participation has been found to improve both robustness and accuracy (Guo et al., 2023). Moreover, recent work demonstrates that mitigating overfocusing yields smoother, more interpretable attention maps and stronger performance on dense prediction and object discovery tasks (Darcet et al., 2024). One approach to mitigate overfocusing is to transform the row-stochastic attention matrix into a doubly-stochastic one, for example, using the Sinkhorn algorithm (Sinkhorn, 1964). While effective in enhancing robustness, this approach incurs even greater computational cost than standard quadratic attention, since Sinkhorn iterations are applied on top of the full quadratic attention map. More recently, Shahbazi et al. (2025) exploited the connection between transport plans (Peyré & Cuturi, 2019) and doubly-stochastic attention matrices and, building on advances in efficient transport plan computation (Liu et al., 2025), introduced ESPFormer—a more computationally efficient formulation of doubly-stochastic attention. Nevertheless, ESPFormer remains quadratic during training due to its reliance on soft-sorting and exhibits super-linear complexity at inference.
@@
    We introduce LOTFormer, a mathematically rigorous mechanism for computing doubly-stochastic attention with linear complexity in sequence length.
@@
    We demonstrate that LOTFormer achieves state-of-the-art performance on the Long Range Arena (LRA) benchmark.
@@
    We show that LOTFormer achieves state-of-the-art performance on ImageNet-1K compared to other linear attention methods.
@@ 2 Related Work
2.1 Linear Attention Mechanisms
The original Transformer architecture relies on the quadratic-time softmax attention, which quickly becomes prohibitive for long sequences (Vaswani et al., 2017). To mitigate this issue, a large body of work has introduced linear attention mechanisms that approximate or restructure the attention computation. Early efforts include Local Attention (Aguilera-Martos et al., 2024) and Longformer (Beltagy et al., 2020), which restrict interactions to a local context window to reduce complexity. More structured sparsity patterns are adopted by Reformer (Kitaev et al., 2020) and BigBird (Zaheer et al., 2020), which employ locality-sensitive hashing and block-sparse attention, respectively.
@@ 2.1 Linear Attention Mechanisms
A second family of approaches replaces the softmax kernel with low-rank or kernelized approximations. Linear Transformer (Katharopoulos et al., 2020) and Performer (Choromanski et al., 2020) leverage kernel feature maps to re-express attention as a dot-product in a lower-dimensional space, reducing complexity to linear in sequence length. Linformer (Wang et al., 2020) introduces low-rank projections of the key–value matrices, while Nyströmformer (Xiong et al., 2021) applies Nyström approximations for kernel matrices. Similarly, Kernelized Attention (Luo et al., 2021), Cosformer (Qin et al., 2022), and Skyformer (Chen et al., 2021) extend this principle with improved kernel feature maps, orthogonal projections, or structured bases. Synthesizer Tay et al. (2021) departs from query–key interactions altogether by learning synthetic attention weights. Finally, Informer (Zhou et al., 2021) and Hedgehog (Zhang et al., 2024) address time-series and structured sequence modeling by combining efficient approximations with task-specific inductive biases. Together, these works establish the central paradigm of linear attention: replacing dense quadratic interactions with kernel, low-rank, or sparse mechanisms that preserve expressivity while improving scalability.
2.2 Doubly-Stochastic Attention
@@
While linear attention addresses efficiency, another line of research has focused on the stochastic structure of attention matrices. The softmax operation enforces row-stochasticity, but not column normalization, leading to asymmetric attention maps. SinkFormer (Sander et al., 2021) addressed this by enforcing doubly-stochastic (DS) constraints via iterative Sinkhorn normalization, thereby producing balanced transport-like couplings between queries and keys. Building on this, ESPFormer (Shahbazi et al., 2025) leveraged expected sliced transport plans with annealed temperature schedules, achieving both theoretical guarantees and empirical improvements over SinkFormer by bridging soft and hard stochastic maps.
More recently, Quantum DS Attention (Born et al., 2025) has extended this framework into quantum-inspired settings, exploiting properties of quantum stochastic matrices and unitary constraints to model richer forms of normalization while preserving tractability. These approaches highlight that double stochasticity is not merely a regularization tool but a principled rethinking of attention as a transport problem, ensuring symmetry, stability, and interpretability in the learned couplings.
2.3 Position of LOTFormer
Our work, LOTFormer, sits at the intersection of these two threads. On the one hand, it inherits the efficiency benefits of linear attention by reformulating the attention kernel in terms of linearizable operations. On the other hand, it extends the doubly stochastic paradigm by embedding transport-inspired constraints into a linearized setting, thereby enabling scalable doubly stochastic attention with provable structure and practical efficiency. By unifying these perspectives, LOTFormer opens a new direction for scalable and structured attention that maintains both computational tractability and principled normalization.
@@ 3 Method
Setup.
Let $\{x_{i}\}_{i=1}^{n}\subset\mathbb{R}^{d_{\mathrm{in}}}$ be the set of $n$ input tokens to the attention block and
$q_{i}\;=\;W_{Q}x_{i}\in\mathbb{R}^{d_{k}},\qquad k_{i}\;=\;W_{K}x_{i}\in\mathbb{R}^{d_{k}},\qquad v_{i}\;=\;W_{V}x_{i}\in\mathbb{R}^{d_{v}},$
@@
Empirical measures and pivot.
In line with earlier approaches that establish a connection between attention mechanisms and optimal transport (OT) (Sander et al., 2022; Shahbazi et al., 2025), we represent queries and keys as empirical measures, namely,
$\mu=\sum_{i=1}^{n}\mathrm{p}^{1}_{i}\delta_{q_{i}},\qquad\nu\;=\;\sum_{j=1}^{n}\mathrm{p}^{2}_{j}\,\delta_{k_{j}},\qquad\text{(typically }\mathrm{p}_{i}^{1}=\mathrm{p}^{2}_{j}=\tfrac{1}{n}\text{)},$
and introduce a learnable pivot measure with support on $r\ll n$ points:
$\sigma\;=\;\sum_{t=1}^{r}\mathrm{p}^{0}_{t}\,\delta_{z_{t}},\qquad Z=\{z_{t}\}_{t=1}^{r}\subset\mathbb{R}^{d_{k}},\;\;\mathrm{p}^{0}_{t}>0,\;\sum_{t=1}^{r}\mathrm{p}^{0}_{t}=1,\;\;r\ll n.$
Conditional OT $\mu\to\sigma\to\nu$.
An alternative but fundamental perspective on attention is that the entropically regularized transport plan between $\mu$ and $\nu$—that is, their joint measure—can be interpreted as a doubly stochastic attention matrix linking queries and keys. Our approach builds on this view by introducing a conditional transport plan, defined relative to a pivot measure, to construct such a coupling. By design, this formulation yields an attention matrix that is both low-rank and linearly decomposable.
Here we define the entropic OT problems from $\mu$ to $\sigma$, and from $\sigma$ to $\nu$. Let us use the dot-product similarity $c(q,k)=q^{\top}k$ and define
@@
$\Gamma^{(1)}$ $\in\arg\max_{\Gamma\in U(\mathrm{p}^{0},\mathrm{p}^{1})}\;\;\langle C^{(1)},\Gamma\rangle\;+\;\varepsilon\,H(\Gamma),$ (3.1)
$\Gamma^{(2)}$ $\in\arg\max_{\Gamma\in U(\mathrm{p}^{0},\mathrm{p}^{2})}\;\;\langle C^{(2)},\Gamma\rangle\;+\;\varepsilon\,H(\Gamma),$ (3.2)
@@
Glued coupling and doubly-stochastic attention.
Define the glued (composed) coupling
$\Gamma\;=\;(\Gamma^{(1)})^{\top}\,\mathrm{Diag}(\sigma)^{-1}\,\Gamma^{(2)}\;\in\;\mathbb{R}^{n\times n}.$ (3.3)
@@
$A\;=\;\,\Gamma\;=\;\,(\Gamma^{(1)})^{\top}\,\mathrm{Diag}(\sigma)^{-1}\,\Gamma^{(2)},\qquad A\mathbf{1}=\mathbf{1},\;\;\mathbf{1}^{\top}A=\mathbf{1}^{\top},$
@@
$\mathrm{LOTAttn}(Q,K,V)\;=\;A\,V\;=\;\,(\Gamma^{(1)})^{\top}\Big(\mathrm{Diag}(\sigma)^{-1}\,(\Gamma^{(2)}V)\Big).$ (3.4)
Rank and complexity.
Write
@@
hence $\mathrm{rank}(A)\leq r$ (Note, in default setting $r\leq n$, we can show $\mathrm{rank}(A)=r$). We may avoid forming $A$ itself by using its factorized form, first computing $Y=\Gamma^{(2)}V\in\mathbb{R}^{r\times d_{v}}$, then $Z^{\prime}=\mathrm{Diag}(\sigma)^{-1}Y$, and finally $O=\Gamma^{(1)}Z^{\prime}\in\mathbb{R}^{n\times d_{v}}$. Each Sinkhorn iteration uses matrix–vector scalings with kernels $K^{(1)}=\exp(QZ^{\top}/\varepsilon)$ and $K^{(2)}=\exp(ZK^{\top}/\varepsilon)$, costing $O(nr)$ per iteration (plus $O(nd_{k}r)$ to form $QZ^{\top}$ and $ZK^{\top}$ once per update). For fixed $r\ll n$ and modest iteration count $T$, the per-head complexity is $O\!\big(nd_{k}r+T\,nr\big)$, i.e., linear in $n$.
Learning.
The pivot locations $Z=\{z_{t}\}$, masses $\sigma\in\Delta^{r-1}$, and projections $(W_{Q},W_{K},W_{V})$ are trained end-to-end by backpropagating through the shared-$\varepsilon$ Sinkhorn problems (3.1)–(3.2) and the glued composition. Note that for multi-head attention, each head has its own $(Z,\sigma)$.
Why not the optimal Low-Rank OT?
Our construction produces an attention matrix that is, by design, low-rank due to the introduction of the pivot measure. However, this should not be conflated with the optimal low-rank transport plan studied in prior work Scetbon & Cuturi (2022); Scetbon et al. (2021). In Low-Rank OT, the support of the low-rank factors is optimized directly to minimize the transportation cost, yielding a low-rank approximation that is provably optimal in that restricted class. By contrast, in our setting, the pivot measure is learned end-to-end as part of the attention mechanism, but not optimized explicitly for transport cost minimization. This distinction naturally raises the question of why not simply employ Low-Rank OT instead? The key reason is computational: computing Low-Rank OT still requires access to the full $n\times n$ cost matrix between $\mu$ and $\nu$ as input, which entails quadratic time and memory. This completely undermines the goal of linear-time attention. In contrast, our conditional transport construction never forms the full cost matrix, but instead only evaluates $QZ^{\top}$ and $ZK^{\top}$, each of size $n\times r$ with $r\ll n$, achieving linear complexity in $n$ while preserving double-stochasticity.
4 Experiments
@@ 4.1 Runtime and Wall-Clock Analysis
Figure 3 reports forward-pass runtime (ms/iteration) versus sequence length on a log–log scale for $N\in\{2^{9:17}\}$. The left panel compares quadratic methods to LOTFormer; the right panel compares linear-time baselines to the same LOTFormer curves. As expected, Softmax and the DS variants exhibit near–quadratic growth in $N$, whereas LOTFormer and the linear baselines grow roughly linearly. Across LOTFormer ranks, lower $r$ yields faster curves and higher $r$ yields slower curves, illustrating a clear rank–time trade-off.
Figure 3: Runtime scaling with sequence length. Forward-pass runtime (ms/iteration) vs. NN on a log–log scale for quadratic methods (left) and linear methods (right). LOTFormer is shown at ranks $r\!\in\!\{16,32,64,128,256\}$ in both panels (solid for r=64r{=}64, dashed otherwise). Points denote measured values at $N\!\in\!\{2^{9:17}\}$;

### v2

# LOTFormer: Doubly Stochastic Linear Attention via Low Rank Optimal Transport

@@ Abstract
Transformers have proven highly effective across modalities, but standard softmax attention scales quadratically with sequence length, limiting long context modeling. Linear attention mitigates this by approximating attention with kernel feature maps, yet most attention mechanisms remain row normalized and can over concentrate mass on a few tokens, harming robustness and information flow. Doubly stochastic attention counteracts this by balancing token participation across both rows and columns, but existing approaches often add significant overhead. We propose LOTFormer, a linear time doubly stochastic attention mechanism derived from an optimal transport view of attention as a coupling between query and key measures. LOTFormer enforces a low rank transport plan by conditioning on a learnable pivot measure with small support. We solve two entropic transport problems, queries to pivot and pivot to keys, and compose them into a conditional coupling that is provably doubly stochastic, has rank at most $r\ll n$, and applies to values in $O(nr)$ time without forming the full $n\times n$ matrix. The pivot locations and masses are learned end-to-end. Across vision and text benchmarks, LOTFormer delivers strong accuracy efficiency tradeoffs when plugged into standard backbones including Swin, DeiT, and BERT.
Machine Learning, ICML
@@ 1 Introduction
On the other hand, prior work has shown that the row-stochastic nature of attention matrices often leads to attention concentrating disproportionately on a few tokens, which can hinder effective information flow (Sander et al., 2022; Shahbazi et al., 2025). In the case of vision Transformers, this phenomenon—referred to as token overfocusing—has been observed to degrade performance, and promoting broader token participation has been found to improve both robustness and accuracy (Guo et al., 2023). Moreover, recent work demonstrates that mitigating overfocusing yields smoother, more interpretable attention maps and stronger performance on dense prediction and object discovery tasks (Darcet et al., 2024). One approach to mitigate overfocusing is to transform the row-stochastic attention matrix into a doubly-stochastic one, for example, using the Sinkhorn algorithm (Sinkhorn, 1964). While effective in enhancing robustness, this approach incurs even greater computational cost than standard quadratic attention, since Sinkhorn iterations are applied on top of the full quadratic attention map. More recently, Shahbazi et al. (2025) exploited the connection between transport plans (Peyré and Cuturi, 2019) and doubly-stochastic attention matrices and, building on advances in efficient transport plan computation (Liu et al., 2025), introduced ESPFormer—a more computationally efficient formulation of doubly-stochastic attention. Nevertheless, ESPFormer remains quadratic during training due to its reliance on soft-sorting and exhibits super-linear complexity at inference.
@@
    We propose LOTFormer, a linear time attention operator that realizes a doubly stochastic attention matrix through a low rank optimal transport factorization via a learnable pivot measure, enabling DS normalization without forming an $n\times n$ attention map.
@@
    We introduce a controlled plug and play evaluation protocol for bidirectional encoder self attention, swapping only the attention modules in trained checkpoints while keeping all other parameters fixed.
@@
    We demonstrate consistent ImageNet 1K gains on DeiT Tiny and Small, PVT Tiny, and Swin Tiny, matching or exceeding strong linear attention and doubly stochastic baselines under matched training and evaluation settings.
@@ 2 Related Work
Linear Attention Mechanisms. The original Transformer relies on quadratic time softmax attention, which becomes prohibitive as sequence length grows (Vaswani et al., 2017). To address this, a broad literature has developed linear attention methods that approximate or restructure attention to reduce computational and memory costs. Early approaches include Local Attention (Aguilera-Martos et al., 2024) and Longformer (Beltagy et al., 2020), which restrict interactions to local windows. More structured sparsity is explored by Reformer (Kitaev et al., 2020) and BigBird (Zaheer et al., 2020), which use locality sensitive hashing and block sparse patterns, respectively. A second family replaces the softmax kernel with low rank or kernelized approximations. Linear Transformer (Katharopoulos et al., 2020) and Performer (Choromanski et al., 2020) use kernel feature maps to rewrite attention so that computation scales linearly with sequence length. Linformer (Wang et al., 2020) introduces low rank projections of key and value representations, while Nyströmformer (Xiong et al., 2021) applies Nyström approximations to kernel matrices. Related directions include Kernelized Attention (Luo et al., 2021), Cosformer (Qin et al., 2022), and Skyformer (Chen et al., 2021), which improve feature maps, projections, or bases. Synthesizer (Tay et al., 2021) departs from explicit query key interactions by learning synthetic attention weights. Informer (Zhou et al., 2021) and Hedgehog (Zhang et al., 2024b) target time series and structured settings, combining efficient approximations with task specific inductive biases. Collectively, these works replace dense quadratic interactions with kernel, low rank, or sparse mechanisms that aim to preserve expressivity while improving scalability.
@@
Doubly Stochastic Attention. Beyond efficiency, another line of work focuses on the stochastic structure of attention matrices. Softmax enforces row stochasticity but does not normalize columns, which can yield imbalanced attention where a small set of tokens receives disproportionate mass. SinkFormer (Sander et al., 2021) enforces doubly stochastic constraints via iterative Sinkhorn normalization, producing balanced transport like couplings between queries and keys. ESPFormer (Shahbazi et al., 2025) leverages expected sliced transport plans with annealed temperature schedules, aiming to improve both stability and performance by bridging soft and hard assignments.
More recently, Quantum DS Attention (Born et al., 2025) extends doubly stochastic constructions using quantum inspired parameterizations, exploiting structure from quantum stochastic matrices and unitary constraints while maintaining tractability. We view this as concurrent work that appeared during our study, and therefore we do not include a direct comparison against it in our experiments. Overall, these methods suggest that doubly stochasticity is more than a heuristic regularizer: it reframes attention as a transport problem with stronger normalization that can improve stability and interpretability of the learned couplings.
Position of LOTFormer. LOTFormer sits at the intersection of efficient and structured attention. It adopts an optimal transport view of attention, but applies it in linear time by forming a low rank coupling through a learnable pivot measure. This construction yields a doubly stochastic attention operator with explicit structure, while avoiding iterative normalization over full $n\times n$ maps. Moreover, the pivot factorization makes the transport structure transparent and interpretable, since interactions are mediated through a small set of representative support points. As a result, LOTFormer can be seen as a principled bridge between doubly stochastic normalizations and practical linear attentions.
@@ 3 Method
We follow the notation in Appendix A for optimal transport and doubly stochastic attention.
@@
Entropic Optimal Transport. Optimal transport (OT) constructs a nonnegative coupling matrix that links two discrete probability vectors while matching their prescribed marginals. Concretely, given $p\in\Delta^{n-1}$, $q\in\Delta^{m-1}$, and a similarity matrix $C\in\mathbb{R}^{n\times m}$, OT searches over the transportation polytope
$U(p,q)\;=\;\{\Gamma\in\mathbb{R}^{n\times m}_{+}:\ \Gamma\mathbf{1}=p,\ \Gamma^{\top}\mathbf{1}=q\},$
and selects a coupling that favors large similarity. Entropic regularization augments this selection with the entropy
$H(\Gamma)\;=\;-\sum_{i,j}\Gamma_{ij}\big(\log\Gamma_{ij}-1\big),$
which encourages diffuse couplings and yields a smooth, stable mapping from similarities to couplings. We will use this template as a normalization primitive, and later instantiate it with dot product similarities.
OT as structured normalization. From the attention viewpoint, an OT coupling turns raw pairwise similarities into weights whose total mass allocation across rows and columns is controlled by the chosen marginals. This provides a principled alternative to purely row normalized attention and motivates constructing attention weights by solving OT problems induced by learned similarities.
Setup. Let $\{x_{i}\}_{i=1}^{n}\subset\mathbb{R}^{d_{\mathrm{in}}}$ be the set of $n$ input tokens to the attention block and
$q_{i}=W_{Q}x_{i}\in\mathbb{R}^{d_{k}},k_{i}=W_{K}x_{i}\in\mathbb{R}^{d_{k}},v_{i}=W_{V}x_{i}\in\mathbb{R}^{d_{v}}.$
@@
Empirical measures and pivot. In line with earlier approaches that establish a connection between attention mechanisms and optimal transport (OT) (Sander et al., 2022; Shahbazi et al., 2025), we represent queries and keys as empirical measures, namely,
$\mu$ $=\sum_{i=1}^{n}\mathrm{p}^{1}_{i}\,\delta_{q_{i}},\qquad\nu=\sum_{j=1}^{n}\mathrm{p}^{2}_{j}\,\delta_{k_{j}},$
$\text{typically }\mathrm{p}^{1}_{i}=\mathrm{p}^{2}_{j}=\tfrac{1}{n}$. We also introduce a learnable pivot measure with support on $r\ll n$ points:
$\sigma$ $=\sum_{t=1}^{r}\mathrm{p}^{0}_{t}\,\delta_{z_{t}},\qquad Z=\{z_{t}\}_{t=1}^{r}\subset\mathbb{R}^{d_{k}},$
$\mathrm{p}^{0}_{t}>0,\quad\sum_{t=1}^{r}\mathrm{p}^{0}_{t}=1,\quad r\ll n.$
Conditional OT $\mu\to\sigma\to\nu$. An alternative but fundamental perspective on attention is that the entropically regularized transport plan between $\mu$ and $\nu$—that is, their joint measure—can be interpreted as a doubly stochastic attention matrix linking queries and keys. Our approach builds on this view by introducing a conditional transport plan, defined relative to a pivot measure, to construct such a coupling. By design, this formulation yields an attention matrix that is both low-rank and linearly decomposable.
Here we define the entropic OT problems from $\mu$ to $\sigma$, and from $\sigma$ to $\nu$. Setting similarity to $c(q,k)=q^{\top}k$ we define
@@
$\Gamma^{(1)}$ $\in\arg\max_{\Gamma\in U(\mathrm{p}^{0},\mathrm{p}^{1})}\;\;\langle C^{(1)},\Gamma\rangle\;+\;\varepsilon\,H(\Gamma),$ (1)
$\Gamma^{(2)}$ $\in\arg\max_{\Gamma\in U(\mathrm{p}^{0},\mathrm{p}^{2})}\;\;\langle C^{(2)},\Gamma\rangle\;+\;\varepsilon\,H(\Gamma),$ (2)
@@
Glued coupling and doubly-stochastic attention. Define the glued (composed) coupling
$\Gamma\;=\;(\Gamma^{(1)})^{\top}\,\mathrm{Diag}(\sigma)^{-1}\,\Gamma^{(2)}\;\in\;\mathbb{R}^{n\times n}.$ (3)
@@
$A\;=\;\Gamma$ $=(\Gamma^{(1)})^{\top}\,\mathrm{Diag}(\sigma)^{-1}\,\Gamma^{(2)},$
$s.t.\quad A\mathbf{1}$ $=\mathbf{1},\qquad\mathbf{1}^{\top}A=\mathbf{1}^{\top}.$
@@
$\mathrm{LOTAttn}(Q,K,V)=AV$ (4)
$=(\Gamma^{(1)})^{\top}\!\Big(\mathrm{Diag}(\sigma)^{-1}\,(\Gamma^{(2)}V)\Big).$
Rank and complexity. Write
@@
hence $\mathrm{rank}(A)\leq r$ (Note, in default setting $r\leq n$, we can show $\mathrm{rank}(A)=r$). We may avoid forming $A$ itself by using its factorized form, by first computing $Y=\Gamma^{(2)}V\in\mathbb{R}^{r\times d_{v}}$, then $Z^{\prime}=\mathrm{Diag}(\sigma)^{-1}Y$, and finally $O=\Gamma^{(1)}Z^{\prime}\in\mathbb{R}^{n\times d_{v}}$. Each Sinkhorn iteration uses matrix–vector scalings with kernels $K^{(1)}=\exp(QZ^{\top}/\varepsilon)$ and $K^{(2)}=\exp(ZK^{\top}/\varepsilon)$, costing $O(nr)$ per iteration (plus $O(nd_{k}r)$ to form $QZ^{\top}$ and $ZK^{\top}$ once per update). For fixed $r\ll n$ and modest iteration count $T$, the per-head complexity is $O\!\big(nd_{k}r+T\,nr\big)$, i.e., linear in $n$.
Learning. The pivot locations $Z=\{z_{t}\}$, masses $\sigma\in\Delta^{r-1}$, and projections $(W_{Q},W_{K},W_{V})$ are trained end-to-end by backpropagating through the shared-$\varepsilon$ Sinkhorn problems (1)–(2) and the glued composition. Note that for multi-head attention, each head has its own $(Z,\sigma)$.
For completeness, we provide a short mathematical explanation of why doubly stochastic constraints are incompatible with standard causal masking except for degenerate solutions. See Appendix  F.
Why not the optimal Low-Rank OT? Our construction produces an attention matrix that is, by design, low-rank due to the introduction of the pivot measure. However, this should not be conflated with the optimal low-rank transport plan studied in prior work (Scetbon and Cuturi, 2022; Scetbon et al., 2021). In Low-Rank OT, the support of the low-rank factors is optimized directly to minimize the transportation cost, yielding a low-rank approximation that is provably optimal in that restricted class. By contrast, in our setting, the pivot measure is learned end-to-end as part of the attention mechanism, but not optimized explicitly for transport cost minimization. This distinction naturally raises the question of why not simply employ Low-Rank OT instead? The key reason is computational: computing Low-Rank OT still requires access to the full $n\times n$ cost matrix between $\mu$ and $\nu$ as input, which entails quadratic time and memory. This completely undermines the goal of linear-time attention. In contrast, our conditional transport construction never forms the full cost matrix, but instead only evaluates $QZ^{\top}$ and $ZK^{\top}$, each of size $n\times r$ with $r\ll n$, achieving linear complexity in $n$ while preserving double-stochasticity.
4 Experiments11footnotemark: 1
22footnotetext: Our code and implementation can be found here.
@@ 4.1 Runtime and Wall-Clock Analysis
Figure 3 reports forward-pass runtime (ms/iteration) versus sequence length on a log–log scale for $N\in\{2^{9:17}\}$. The left panel compares quadratic methods to LOTFormer; the right panel compares linear-time baselines to the same LOTFormer curves. As expected, Softmax and the DS variants exhibit near–quadratic growth in $N$, whereas LOTFormer and the linear baselines grow roughly linearly. Across LOTFormer ranks, lower $r$ yields faster curves and higher $r$ yields slower curves, illustrating a clear rank–time trade-off. We include an appendix note discussing how LOTFormer differs from FlashAttention at the kernel level, and why our current implementation is an unfused PyTorch baseline rather than a fused attention kernel. See Appendix  E for the scope disclaimer.
Figure 3: Runtime scaling with sequence length. Forward-pass runtime (ms/iteration) vs. NN on a log–log scale for quadratic methods (left) and linear methods (right). LOTFormer is shown at ranks $r\!\in\!\{16,32,64,128,256\}$ in both panels (solid for r=64r{=}64, dashed otherwise). Points denote measured values at $N\!\in\!\{2^{9:17}\}$.
4.2 ImageNet-1K
Next, we evaluate LOTFormer on ImageNet 1K at $224^{2}$ input resolution with DeiT Tiny (Touvron et al., 2021), DeiT Small, PVT Tiny (Wang et al., 2021), and Swin Tiny (Liu et al., 2021), and compare against recent efficient attention variants (Table 1). LOTFormer achieves the best Top 1 within the DeiT Tiny, DeiT Small, and PVT Tiny blocks, improving over the corresponding Softmax baselines from $72.2$ to $74.8$, from $79.8$ to $80.1$, and from $75.1$ to $79.6$, respectively, with comparable parameter counts and FLOPs. On Swin Tiny, LOTFormer attains $81.8$ Top 1, placing it as the top performer in most blocks and within the top three in the remaining block at the same resolution.
[CLS] under DS attention. A practical issue when using doubly stochastic (DS) attention in ViTs is the [CLS] token: it must pool information from all image tokens, while DS constraints can, in principle, over-concentrate attention and weaken global aggregation. In the extreme case (vanishing entropy regularization and $r=n$), a DS attention matrix may approach a permutation, so each query attends to a single key, and [CLS] would aggregate from only one token. Although entropy regularization and low-rank structure mitigate this in practice, the tension remains conceptually. To preserve the global pooling role of [CLS], we keep a standard softmax aggregator for the [CLS] row and apply DS constraints only among the image tokens:
$\alpha_{\mathrm{cls},j}=\frac{\exp\!\big(\beta\,\langle q_{\mathrm{cls}},k_{j}\rangle\big)}{\sum_{l}\exp\!\big(\beta\,\langle q_{\mathrm{cls}},k_{l}\rangle\big)},\qquad\beta>0,$ (5)
while all non-[CLS] rows follow the DS attention rule.
[CLS] Polarization. Following (Meng et al., 2025), we split a scalar $x$ into its positive and negative components via $x^{+}=\mathrm{ReLU}(x)$ and $x^{-}=\mathrm{ReLU}(-x)$. In LOTFormer, we optionally restrict this operation to the [CLS] token, so that only [CLS]-to-token (and token-to-[CLS]) interactions are polarized while all other token-to-token interactions remain unchanged, by defining:
$\tilde{s}_{\mathrm{cls},j}$ $=\Big((q_{\mathrm{cls}}^{+})^{\top}k_{j}^{+}+(q_{\mathrm{cls}}^{-})^{\top}k_{j}^{-}\Big)^{p_{s}}$ (6)
$\quad+\Big((q_{\mathrm{cls}}^{+})^{\top}k_{j}^{-}+(q_{\mathrm{cls}}^{-})^{\top}k_{j}^{+}\Big)^{p_{o}},$
followed by $\alpha_{\mathrm{cls},j}=\mathrm{softmax}_{j}(\tilde{s}_{\mathrm{cls},j})$. By contrast, PolaFormer applies the same polarized logit construction to all rows, followed by per-row softmax normalization. In our setting, only the [CLS] row uses softmax (with optional polarization), while the remaining rows remain DS.
Reporting protocol. Table 1 reports the best configuration per method at $224^{2}$ resolution for each backbone family. Several efficient attention baselines, most notably PolaFormer, include a depthwise convolution (DWC) token mixing layer as part of their reported architecture. To ensure fair comparisons, we run LOTFormer under the same setting when comparing against such methods, and we additionally report matched results without DWC. Table 3 isolates the effect of DWC under otherwise identical training and evaluation settings, together with ablations on the [CLS] aggregator and [CLS] only polarization. Across backbones, the dominant gain for doubly stochastic style attention comes from enabling a dedicated [CLS] softmax aggregator, while [CLS] only polarization provides a smaller but consistent improvement.
[…]
Table 1: ImageNet 1K results at $224^{2}$ input resolution across four backbones. We report parameter count, FLOPs, and Top 1 accuracy. Best Top 1 within each backbone block is in bold.
Method Input Params FLOPs Top 1
DeiT Tiny
[…]
LOTFormer $224^{2}$ 5.8M 1.2G 74.8
DeiT Small
[…]
LOTFormer $24^{2} 22$.3M 4.8G 80.1
PVT Tiny
[…]
LOTFormer $224^{2}$ 12.2M 2.1G 79.6
Swin Tiny
[…]
LOTFormer $224^{2}$ 28.0M 4.5G 81.8